determine the amplitude and period of each function graph each function from -2pi to 2pi y=-1/4 cos x/2
Answer by stanbon(75887)

determine the amplitude and period of each function graph each function from -2pi to 2pi y=-1/4 cos x/2
--------------
amp = |-1/4| = 1/4
---
period = (2pi)/(1/2) = 4pi
=============================
Cheers,
Stan H.